Solve the following
z(p^2+q^2)+px+qy=0
1
Expert's answer
2020-05-19T16:09:50-0400

Given "(p^2+q^2)z+px+qy = 0" _______________________(1)

Let "f = (p^2+q^2)z+px+qy"

Using Charpit's Method:

"\frac{dx}{-\frac{\partial f}{\partial p}} = \frac{dy}{-\frac{\partial f}{\partial q}} = \frac{dz}{-p\frac{\partial f}{\partial p}-q\frac{\partial f}{\partial q}} = \frac{dp}{\frac{\partial f}{\partial x}+p\frac{\partial f}{\partial z}} = \frac{dq}{\frac{\partial f}{\partial y}+q\frac{\partial f}{\partial z}}"

From last two factor, we get

"\frac{dp}{p+p(p^2+q^2)} = \frac{dq}{q+q(p^2+q^2)}" "\implies \frac{dp}{p} = \frac{dq}{q}"

So by integrating, we get "log(p) = log(q) - log(a) \implies q = ap".

By putting value of "q" in (1), we get

"p^2(1+a^2)z + px+apy = 0 \\ 
\implies p (1+a^2)z + (x+ay) = 0 \\
\implies p = -\frac{x+ay}{(1+a^2)z} \ and \ q = -\frac{a(x+ay)}{(1+a^2)z}".

Finally, we have to solve: "dz = pdx+qdy"

"dz = -\frac{x+ay}{(1+a^2)z} dx -\frac{a(x+ay)}{(1+a^2)z} dy"

"\implies (1+a^2)zdz = -xdx - a(ydx+xdy) -a^2 ydy"

So by integrating, we get

"(1+a^2)\frac{z^2}{2} = -\frac{x^2}{2} - axy -a^2 \frac{y^2}{2}"

"\implies (1+a^2)z^2 = -x^2 - 2axy -a^2 y^2" is the solution of given differential equation.